FOURIER TRANSFORM EMISSION SPECTROSCOPY OF NEW ELECTRONIC TRANSITIONS OF TiCl and TiF

\underline{R.S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; P. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

The emission spectra of TiCl and TiF have been investigated at high resolution in the red and near infrared regions. TiF and TiCl were made in a microwave discharge lamp using a mixture of TiCl_{4} or TiF_{4} vapor and He, and the spectra were recorded using a Fourier transform spectrometer. Three new TiCl bands with origins near 6938.9, 6900.2 and 6861.7 cm^{−1} have been assigned as the 0-0, 1-1 and 2-2 bands of a new ^{2}Σ −^{2} Σ transition. This assignment is supported by our recent ab initio calculations on TiCl and ZrCl [Ram et al., J. Chem. Phys. 114, 3977-3987 (2001)]. The new TiF bands are located in the 12000-14000 cm^{−1} spectral region with the 0-0 band near 13232 cm^{−1}. These bands consist of four sub-bands, ^{4}Δ_{1/2} −^{4} Φ_{3/2}, ^{4}Δ_{3/2} −^{4} Φ_{5/2}, ^{4}Δ_{5/2} −^{4} Φ_{7/2} and ^{4}Δ_{7/2} −^{4} Φ_{9/2}, of a ^{4}Δ − X^{4}Φ transition. A rotational analysis of the 0-0 and 0-1 bands has been obtained and the spectroscopic constants have been extracted.